# Object Fidelity Diffusion for Remote Sensing Image Generation

### v2

@@ Abstract
High-precision controllable remote sensing image generation is both meaningful and challenging. Existing diffusion models often produce low-fidelity images due to their inability to adequately capture morphological details, which may affect the robustness and reliability of object detection models. To enhance the accuracy and fidelity of generated objects in remote sensing, this paper proposes Object Fidelity Diffusion (OF-Diff), which effectively improves the fidelity of generated objects. Specifically, we are the first to extract the prior shapes of objects based on the layout for diffusion models in remote sensing. Then, we introduce a dual-branch diffusion model with diffusion consistency loss, which can generate high-fidelity remote sensing images without providing real images during the sampling phase. Furthermore, we introduce DDPO to fine-tune the diffusion process, making the generated remote sensing images more diverse and semantically consistent. Comprehensive experiments demonstrate that OF-Diff outperforms state-of-the-art methods in the remote sensing across key quality metrics. Notably, the performance of several polymorphic and small object classes shows significant improvement. For instance, the mAP increases by 8.3%, 7.7%, and 4.0% for airplanes, ships, and vehicles, respectively.
@@
Figure 1: Four critical failure modes in the State-of-the-Art (SOTA) method ($C$-Diff): (a) Control Leakage: content spills outside the specified layout; (b) Structural Distortion: resulting in malformed object morphology; (c) Dense Generation Collapse: a loss of control in crowded scenes; and (d) Feature-level Mismatch: a distributional drift from real data, visualized by t-SNE. Our OF-Diff (2nd row) effectively resolves these issues.
Introduction
@@
Synthesizing high-fidelity, spatially-controllable remote sensing (RS) images is a critical frontier for overcoming the data limitations that hinder downstream perception tasks like object detection (Yang et al. 2021; Zhang et al. 2020; Yang et al. 2019). Current RS generation methods, however, typically rely on either ambiguous text prompts (Khanna et al. 2023; Sebaq and ElHelw 2024) or auxiliary conditions like semantic maps (Sebaq and ElHelw 2024; Tang et al. 2024; Gong et al. 2024; Hu et al. 2025; Jia et al. 2025). While visually plausible, such guidance is fundamentally disconnected from the instance-level ground truth, failing to provide the precise control necessary for effective data augmentation.
@@
In contrast, Layout-to-Image (L2I) generation conditioned on object bounding boxes presents a more robust solution for precise spatial control. While this approach has been extensively developed in the natural image domain with methods like LayoutDiffusion (Zheng et al. 2023) treating it as a multi-modal fusion task, GLIGEN (Li et al. 2023) enabling open-world generation with extra controls, and ODGen (Zhu et al. 2024) decoupling objects for higher controllability, its direct application to RS is non-trivial, due to the unique challenges of RS imagery, such as vast backgrounds, arbitrary object orientations, and dense scenes.
@@
In RS layout-to-image generation, existing methods like AeroGen (Tang et al. 2025) and CC-Diff (Zhang et al. 2024) take different approaches. AeroGen, a coarse layout-conditioned model, suffers from limited spatial and shape control. In contrast, instance-level methods like CC-Diff achieve higher controllability and fidelity by referencing real instances, but this creates heavy dependence on the quality and quantity of real data, limiting generalization and flexibility. We summarize common failure modes (see Figure 1), including control leakage, structural distortion, dense generation collapse, and feature-level mismatch.
@@
These deficiencies significantly degrade the performance on object detection tasks, limiting their practical application in intelligent RS interpretation. In this paper, we introduce Object Fidelity Diffusion Model (OF-Diff). It is designed to improve the shape fidelity and layout consistency of object generation in RS images. As shown in Figure 2, the existing L2I methods are mainly divided into two categories. The first is layout-conditioned baseline, as shown in Figure 2(a), like Aerogen and LayoutDiffusion. The second is the method with instance-based module, as shown in Figure 2(b), like CC-Diff. However, such methods require real instances and images as references during the sampling stage in order to generate high-quality synthetic images. In contrast, OF-Diff only needs to extract the shape of the foreground to generate high-fidelity RS images, as shown in Figure 2(c). In addition, it fine-tunes the diffusion with DDPO, effectively enhancing the performance of downstream tasks for the generated images. The results in Figure 2(d) demonstrates the superiority of OF-Diff over other methods. Our contributions are summarized as follows:
@@
  • We introduce OF-Diff, a dual-branch controllable diffusion model with prior shape extraction, which improves generation fidelity while reducing reliance on real images, enhancing practical applicability.
@@
Figure 3: OF-Diff’s overall architecture. (a) During the training phase, the shape features of objects obtained through ESGM and image are feature extracted by Control, and the parameters are updated to the dual-branch stable diffusion. Noisy image $Z_{t}$ in the latent space is processed through shape feature stable diffusion and Mix feature stable diffusion, generating the noise predictions $\varepsilon_{\theta}^{s}$ and $\varepsilon_{\theta}^{m}$. The diffusion consistency loss is calculated by $\varepsilon_{\theta}^{s}$ and $\varepsilon_{\theta}^{m}$, and the denoising loss is calculated by $\varepsilon_{\theta}^{s}$ and $\varepsilon$. (b) During the sampling phase, only label and shape feature stable Diffusion are utilized to generate synthetic images. (c) The pipeline of OF-Diff. DDPO is optional if the generated data is expected to better conform to the distribution of a specific dataset.
Related Work
Advances in Image Generation
@@
Layout-to-Image Generation
Controllable image synthesis primarily includes text-to-image (T2I) and layout-to-image (L2I) generation. While T2I models (Nichol et al. 2022; Ramesh et al. 2022) achieve semantic alignment via textual prompts, L2I methods offer better spatial control. Recent works enhance layout conditioning through layout-as-modality designs (Zheng et al. 2023), gated attention (Li et al. 2023), and instance-wise generation (Wang et al. 2024; Zhou et al. 2024). However, these methods rely solely on coarse layout inputs (e.g., bounding boxes), which lack fine-grained shape information critical for synthesizing morphologically complex objects.
Remote Sensing Image Synthesis
Synthesizing high-fidelity training data is crucial for advancing remote sensing (RS) object detection, a field critical to numerous applications, but often hampered by the scarcity of extensively annotated datasets. Despite its necessity, most generative models for RS imagery, such as DiffusionSat (Khanna et al. 2023) and RSDiff (Sebaq and ElHelw 2024), still rely on coarse semantic guidance. While other approaches leverage diverse control signals (Tang et al. 2024) like OpenStreetMaps (Espinosa and Crowley 2023), they are generally not optimized for the bounding box format central to object detection. This naturally motivates L2I approaches including AeroGen (Tang et al. 2025) and CC-Diff (Zhang et al. 2024), which have improved spatial accuracy and contextual consistency through layout-mask attention and FG/BG dual re-samplers. However, they suffer from limited controllability and heavy reliance on real data.
Method
Preliminary
Diffusion models (Song, Meng, and Ermon 2021) aim to capture the underlying data distribution $p(x)$ by iteratively reconstructing data from a noisy representation that is initially sampled from a standard normal distribution.
Denoising Diffusion Probabilistic Models (Ho, Jain, and Abbeel 2020) parameterize the model as the function $\epsilon_{\theta}(x_{t},t)$ to predict the noise component of the sample $x_{t}$ at any time step $t$. The training objective is to minimize the mean squared error (MSE) loss between the actual noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},t)$:
@@
Stable Diffusion (Rombach et al. 2022a; Qiu et al. 2025) utilizes a pre-trained VQ-VAE (Van Den Oord, Vinyals et al. 2017) to encode images into a lower-dimensional latent space, performing training on the latent representation $z_{0}$. In the context of conditional generation, given a text prompt $c_{t}$ and task-specific conditions $c_{f}$, the diffusion training loss at time step $t$ can be expressed as:
@@
where $\mathcal{L}$ represents the overall learning objective of the complete diffusion model. This objective function is explicitly applied during the fine-tuning of diffusion models in conjunction with ControlNet (Zhang, Rao, and Agrawala 2023).
Architecture of OF-Diff
As illustrated in Figure 3(a), during the training phase of OF-Diff, both the real images and their corresponding labels are needed for diffusion training. First, they are fed into the Enhanced Shape Generation Module (ESGM) for shape extraction, yielding the object mask. Subsequently, the real image and its object mask are fed into ControlNet for feature extraction, resulting in the image feature $c_{i}$ and the label feature $c_{l}$. Second, the input image is first compressed into the latent space features $z_{0}$ by a pre-trained VQ-VAE. Then, it is concatenated with random Gaussian noise $\epsilon$ to form $z_{t}$. After being injected into the stable diffusion encoder blocks, the feature $Z_{t}$ is fed into a dual-branch decoder architecture. The two components of this architecture are respectively designated as the shape feature stable diffusion decoder and the mix feature stable diffusion decoder. Third, a weighted summation of $c_{i}$ and the stop - gradient $c_{l}$ is carried out to $c_{m}$. $c_{m}$ is defined as:
$c_{\operatorname{m}}=\frac{n}{N}\cdot c_{i}+\operatorname{sg}\left[c_{l}\right],$ (3)
where $n$ denotes the current iteration number, and $N$ is the total number of iterations. In order to enable the prediction from the mix feature stable diffusion to serve as an anchor point, guiding the convergence trajectory of the shape stable diffusion and thus improving the morphological fidelity of the image, we adopt a stop-gradient strategy (Chen and He 2021) for $c_{l}$ when calculating $c_{m}$. The loss $\mathcal{L}_{s}$ and $\mathcal{L}_{m}$ of shape feature stable diffusion and mix feature stable diffusion are defined as:
@@
To assist the shape feature stable diffusion in breaking away from local minima with low fidelity and migrating towards regions of high fidelity in the parameter space, we introduce a diffusion consistency loss:
$\mathcal{L}_{c}=\mathbb{E}\left[\left\|\epsilon_{\theta}^{m}-\operatorname{sg}\left[\epsilon_{\theta^{\prime}}^{\operatorname{mix}}\right]\right\|^{2}\right],$ (6)
The predicted noise $\epsilon_{\theta}^{m}$ from the mix feature stable diffusion is more accurate than $\epsilon_{\theta}^{s}$ predicted by the shape feature stable diffusion, owing to the additional image prior control. Nevertheless, during the sampling process, we do not utilize inputs with real images. Consequently, we employ a stop-gradient operation so that the more accurate $\epsilon_{\theta}^{m}$ can serve as an anchor, guiding the convergence trajectory of the masked diffusion towards high-fidelity local minima in the parameter space. Ultimately, the loss employed to train OF-Diff is defined as:
$\mathcal{L}=\mathcal{L}_{s}+\mathcal{L}_{m}+\mathcal{L}_{c},$ (7)
@@
Figure 4: The architecture of Enhanced Shape Generation Module (ESGM).
Enhanced Shape Generation Module
@@
In natural imagery, perspective and scale changes prevent a unique geometric model for most objects. Conversely, remote-sensing objects display quasi-invariant shapes. For instance, courts are rectangular, chimneys and oil tanks circular, and airplanes bilaterally symmetric with a distinct nose and tail. This shape consistency enables the use of masks to impose strong controllability on image synthesis for remote sensing. To better leverage category labels for object shape extraction, we introduce the Enhanced Shape Generation Module (ESGM, see Figure 4). During the training phase, ESGM uses paired images and labels to generate precise object masks. And at sampling time, it employs learned shape priors to synthesize diverse masks of object shape.
@@
For the given image $x_{i}$ and its bounding box $y_{i}^{j}$ corresponding to category $j$ $(j\in[1,N])$, we first utilize the RemoteCLIP (Liu et al. 2024) to generate a textual description of the object enclosed within the bounding box. With this description and the original image $x_{i}$, the RemoteSAM (Yao et al. 2025) then generates the corresponding shape mask $x_{i}^{j}$.
In the shape augmentation phase, we crop $x_{i}^{j}$ using its bounding box $y_{i}^{j}$ to extract a shape mask $x_{i}^{j}$ that contains only the target object. Subsequently, we apply a random rotation to $x_{i}^{j}$ and reposition the rotated mask onto a blank canvas using the original bounding box coordinates $y_{i}^{j}$, resulting in a shape-enhanced mask. During the training phase, ESGM can directly obtain the shape of real images. While sampling, ESGM selects enhanced shapes from a mask pool collected during or beyond training, as in common data-efficient practices where segmentation masks are reused or synthesized with minimal impact on generalization. In our setup, we use masks generated during training.

### v3

@@ Abstract
High-fidelity, controllable remote sensing layout-to-image generation is highly valuable for providing high-quality data for downstream object detection tasks. However, existing methods either rely on additional textual guidance, leading to geometric distortions, or require extra real-image references, limiting practical applicability. To address these challenges, we propose Object Fidelity Diffusion (OF-Diff), which leverages object layouts to extract structural shape priors and employs an online-distillation strategy to integrate complex image features. This allows the model to perform highly controllable, high-fidelity image generation at inference without relying on real-image references. Furthermore, we introduce DDPO to fine-tune the diffusion process, making the generated remote sensing images more diverse and semantically consistent. Comprehensive experiments demonstrate that OF-Diff outperforms state-of-the-art methods in the remote sensing across key quality metrics. Notably, the performance of several polymorphic and small object classes shows significant improvement. For instance, the mAP increases by 8.3%, 7.7%, and 4.0% for airplanes, ships, and vehicles, respectively.
[…]
1 Introduction
@@
Figure 1: Four critical failure modes in the State-of-the-Art (SOTA) method ($C$-Diff): a distributional drift from real data, visualized by t-SNE; and (a) Control Leakage; (b) Structural Distortion; (c) Dense Generation Collapse. Our OF-Diff (2nd row) effectively resolves these issues.
@@
Synthesizing high-fidelity, spatially-controllable remote sensing (RS) images is a critical frontier for overcoming the data limitations that hinder downstream perception tasks like object detection [yang2021r3det, zhang2020well, Yang_2019_ICCV]. Current RS generation methods, however, typically rely on either ambiguous text prompts [khanna2023diffusionsat, sebaq2024rsdiff] or auxiliary conditions like semantic maps [sebaq2024rsdiff, tang2024crs, gong2024crossearth, hu2025earth, jia2025can]. While visually plausible, such guidance is fundamentally disconnected from the instance-level ground truth, failing to provide the precise control necessary for effective data augmentation.
@@
In contrast, Layout-to-Image (L2I) generation conditioned on object bounding boxes offers a more robust solution for precise spatial control. This paradigm has been extensively studied in the natural image domain—LayoutDiffusion [zheng2023layoutdiffusion] treats it as a multi-modal fusion problem, GLIGEN [li2023gligen] enables open-world generation through additional control signals, and ODGen [zhu2024odgen] improves controllability by decoupling objects—yet its direct application to remote sensing (RS) imagery remains non-trivial due to expansive backgrounds, arbitrary object orientations, and densely packed scenes.
@@
In RS layout-to-image generation, existing methods like AeroGen [tang2025aerogen] and CC-Diff [zhang2024cc] take different approaches. AeroGen, a coarse layout-conditioned model, suffers from limited spatial and shape control. In contrast, instance-level methods like CC-Diff achieve higher controllability and fidelity by referencing real instances, but this creates heavy dependence on the quality and quantity of real data, limiting generalization and flexibility.The images generated via CC-Diff diverge more markedly from the real remote sensing data distribution, aligning instead with the style characteristic of the model’s pre-training corpus. We summarize common failure modes (see Figure 1), including control leakage, structural distortion, dense generation collapse and feature-level mismatch.
@@
These deficiencies significantly degrade the performance on object detection tasks, limiting their practical application in intelligent RS interpretation. In this paper, we introduce Object Fidelity Diffusion Model (OF-Diff). It is designed to improve the shape fidelity and layout consistency of object generation in RS images. As shown in Figure 2, the existing L2I methods are mainly divided into two categories. The first is layout-conditioned baseline, as shown in Figure 2(a), like Aerogen and LayoutDiffusion. The second is the method with instance-based module, as shown in Figure 2(b), like CC-Diff. However, such methods require real instances and images as references during the sampling stage in order to generate high-quality synthetic images. In contrast, OF-Diff generates high-fidelity remote-sensing objects using only the foreground shape, and subsequently employs online-distillation to further align the outputs with real images, as shown in Figure 2(c). In addition, it fine-tunes the diffusion with DDPO, effectively enhancing the performance of downstream tasks for the generated images. The results in Figure 2(d) demonstrates the superiority of OF-Diff over other methods. Our contributions are summarized as follows:
@@
  • We introduce OF-Diff, an online-distillation controllable diffusion model with prior shape extraction, which improves generation fidelity while reducing reliance on real images, enhancing practical applicability.
@@
2 Related Work
2.1 Advances in Image Generation
@@
2.2 Layout-to-Image Generation
Controllable image synthesis primarily includes text-to-image (T2I) and layout-to-image (L2I) generation. While T2I models [nichol2022glide, ramesh2022hierarchical] achieve semantic alignment via textual prompts, L2I methods offer better spatial control. Recent works enhance layout conditioning through layout-as-modality designs [zheng2023layoutdiffusion], gated attention [li2023gligen], and instance-wise generation [wang2024instancediffusion, Zhou_2024_CVPR]. However, these methods rely solely on coarse layout inputs (e.g., bounding boxes), which lack fine-grained shape information critical for synthesizing morphologically complex objects.
2.3 Remote Sensing Image Synthesis
Synthesizing high-fidelity training data is crucial for advancing remote sensing (RS) object detection, a field critical to numerous applications, but often hampered by the scarcity of extensively annotated datasets. Despite its necessity, most generative models for RS imagery, such as DiffusionSat [khanna2023diffusionsat] and RSDiff [sebaq2024rsdiff], still rely on coarse semantic guidance. While other approaches leverage diverse control signals [tang2024crs] like OpenStreetMaps [espinosa2023generate], they are generally not optimized for the bounding box format central to object detection. This naturally motivates L2I approaches including AeroGen [tang2025aerogen] and CC-Diff [zhang2024cc], which have improved spatial accuracy and contextual consistency through layout-mask attention and FG/BG dual re-samplers. However, they suffer from limited controllability and heavy reliance on real data.
3 Method
3.1 Preliminary
Diffusion models [song2021denoising] aim to capture the underlying data distribution $p(x)$ by iteratively reconstructing data from a noisy representation that is initially sampled from a standard normal distribution. Denoising Diffusion Probabilistic Models [ho2020denoising] parameterize the model as the function $\epsilon_{\theta}(x_{t},t)$ to predict the noise component of the sample $x_{t}$ at any time step $t$. The training objective is to minimize the mean squared error (MSE) loss between the actual noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},t)$:
@@ 3.1 Preliminary
Stable Diffusion (SD) [Rombach_Blattmann_Lorenz_Esser_Ommer_2022, qiu2025noise] utilizes a pre-trained VQ-VAE [van2017neural] to encode images into a lower-dimensional latent space, performing training on the latent representation $z_{0}$. In the context of conditional generation, given a text prompt $c_{t}$ and task-specific conditions $c_{f}$, the diffusion training loss at time step $t$ can be expressed as:
@@
where $\mathcal{L}$ represents the overall learning objective of the complete diffusion model. This objective function is explicitly applied during the fine-tuning of diffusion models in conjunction with ControlNet [zhang2023adding].
Figure 3: OF-Diff’s overall architecture. (a) During training, object shape features extracted by ESGM and image features are processed by ControlNet, and the resulting information is used to update stable diffusion decoders via online-distillation. (b) During sampling, only the label and the shape feature stable diffusion decoder are used to generate synthetic images. (c) Architecture of the Enhanced Shape Generation Module (ESGM).
3.2 Architecture of OF-Diff
As illustrated in Figure 3(a), the training of OF-Diff requires both real images and their corresponding labels. First, for ControlNet, the real image and its label are processed by the Enhanced Shape Generation Module (ESGM) to extract the object mask. The image and mask are then fed into ControlNet to obtain the image feature $c_{i}$ and the shape feature $c_{s}$. To enrich the structural-only shape prior with richer appearance and contextual cues from the image, we combine them into a mix-feature $c_{m}$, which will later serve as a teacher input in online-distillation. Concretely:
$c_{\operatorname{m}}=\frac{n}{N}\cdot c_{i}+\operatorname{sg}\left[c_{s}\right],$ (3)
where $n$ denotes the current iteration number, and $N$ is the total number of iterations. In order to enable the prediction conditioned on mix-feature to serve as a stable anchor point, to improve the morphological fidelity of the generation, we adopt a stop-gradient strategy [Chen_He_2021] for $c_{s}$ when calculating $c_{m}$.
Second, for Stable Diffusion, the input image is first compressed into latent space features $z_{0}$ by a pre-trained VQ-VAE. Then, it is concatenated with Gaussian noise $\epsilon$ to form $z_{t}$. After passing through the SD encoder blocks, the feature $Z_{t}$ is fed into a dual-decoder architecture. One branch, the shape-feature SD decoder, conditions on $c_{s}$; the other, the mix-feature SD decoder, conditions on $c_{m}$. Their reconstruction losses are defined as $L_{s}$ and $L_{M}$, respectively:
@@ 3.2 Architecture of OF-Diff
Third, for online distillation, the mix-feature SD produces more accurate predictions thanks to its stronger image prior, but needs real images, limiting diversity. In contrast, the shape-feature SD supports arbitrary label control but risks converging to low-fidelity local minima. To reconcile these trade-offs, we propose an online-distillation framework with a consistency loss $L_{c}$:
$\mathcal{L}_{c}=\mathbb{E}\left[\left\|\epsilon_{\theta}^{s}-\operatorname{sg}\left[\epsilon_{\theta^{\prime}}^{\operatorname{m}}\right]\right\|^{2}\right].$ (6)
Here, the prediction $\epsilon_{\theta^{\prime}}^{m}$ from mix-feature SD decoder acts as a stop-gradient teacher signal, serving as an anchor to guide the prediction $\epsilon_{\theta}^{s}$ from shape-feature SD decoder towards high-fidelity optima in parameter space.
The overall training objective is therefore:
$\mathcal{L}=\mathcal{L}_{s}+\mathcal{L}_{m}+\lambda\mathcal{L}_{c},$ (7)
@@
3.3 Enhanced Shape Generation Module
@@
In natural imagery, perspective and scale changes prevent a unique geometric model for most objects. Conversely, remote-sensing objects display quasi-invariant shapes. For instance, courts are rectangular, chimneys and oil tanks circular, and airplanes bilaterally symmetric with a distinct nose and tail. This shape consistency enables the use of masks to impose strong controllability on image synthesis for remote sensing. To better leverage category labels for object shape extraction, we introduce the Enhanced Shape Generation Module (ESGM, see Figure 3(c)). During the training phase, ESGM uses paired images and labels to generate precise object masks. And at sampling time, it employs learned shape priors to synthesize diverse masks of object shape.
@@
For the given image $x_{i}$ and its bounding box $y_{i}^{j}$ corresponding to category $j$ $(j\in[1,N])$, we first utilize the RemoteCLIP [liu2024remoteclip] to generate a textual description of the object enclosed within the bounding box. With this description and the original image $x_{i}$, the RemoteSAM [yao2025remotesam] then generates the corresponding shape masks $x_{i}^{j}$}.
In the shape augmentation phase, each object mask $x_{i}^{j}$ is cropped by its bounding box $y_{i}^{j}$, randomly rotated, and placed back onto a blank canvas to produce a shape-enhanced mask. During training, ESGM uses real image shapes; at sampling, it selects enhanced shapes from a lightweight mask pool collected during or after training. In our experiments, we use masks generated during training.
3.4 DDPO fine-tuning
To enhance the diversity of the distribution of data generated by the fine-tuned model and maintain better consistency with the distribution of real images [schulman2015trust, schulman2017proximal], denoising diffusion policy optimization (DDPO) [black2023training] is applied in the post-training of OF-Diff. DDPO regards the denoising process of the diffusion model as a multi-step Markov decision process (MDP) (for a detailed derivation, please refer to the Appendix 6.1). To optimize the policy $\pi(a_{t}\mid s_{t})$ so as to maximize the cumulative reward $\mathbb{E}_{\tau\sim p(,\cdot\mid\pi)}!\Bigl[\sum_{t=0}^{T}R(\mathbf{s}_{t},\mathbf{a}_{t})\Bigr]$, the gradient $\hat{g}$ is computed as follows:
$\hat{g}=\mathbb{E}\!\biggl[\sum_{t=0}^{T}\frac{p_{\theta}(\mathbf{x}_{t-1}\mid c,\ t,\ \mathbf{x}_{t})}{p_{\theta^{\prime}}(\mathbf{x}_{t-1}\mid c,\ t,\ \mathbf{x}_{t})}\cdot r(\mathbf{x}_{0},c)\cdot\nabla_{\theta}\log p_{\theta}(\mathbf{x}_{t-1}\mid c,\ t,\ \mathbf{x}_{t})\biggr]$ (8)
The reward functions based on K-Nearest Neighbor (KNN) and KL divergence are introduced, respectively, to optimize the diversity of generated data and the distribution consistency between generated data and real data. $\omega$ is the weight parameter, and $\mathbf{x}_{0}^{\prime}$ is the real image in the dataset. Following standard practice, we compute the KNN in the low-dimensional embedding space of CLIP’s image encoder. The implementation details are shown in Appendix 6.1.
4 Experiments
4.1 Experimental Settings
Datasets. DIOR-R [cheng2022anchor], the rotated variant of DIOR [li2020object], contains 20 categories annotated with oriented bounding boxes; we follow the official 1:1:2 split for training/validation/testing. DOTA-v1.0 [Xia_2018_CVPR] includes 15 categories featuring dense scenes and small objects. We crop the images from DOTA to $512\times 512$ following MMRotate [zhou2022mmrotate], discarding those without valid objects. HRSC2016 [hrsc2016] is a high‐resolution ship detection dataset with a multi‐level hierarchical taxonomy. We use the finest‐grained level, consisting of 26 detailed ship categories. The experiments of this dataset are reported in Appendix 6.3 Unless otherwise specified, we train the diffusion model on the trainset. For downstream detection, we use the trainset annotations as layout and mix generated samples with the real trainset, and report evaluation results on the testset.
Implementation Details. We train OF-Diff separately on each dataset (DIOR/DOTA), based on the Stable Diffusion 1.5 [Rombach_2022_CVPR] pretrained model. Only the ControlNet and shape feature SD decoder are fine-tuned, while all other modules remain frozen. The weighting coefficient $\lambda$ of the consistency loss is set to 1, the k value in KNN is set to 50, and the weight $\omega$ of the KL divergence is set to 2. Training is performed using the AdamW optimizer with a learning rate of 1e-5. The global batch size is set to 64, and training runs for 100 epochs.
Benchmark Methods. We compare our method with state-of-the-art L2I generation models for both remote sensing (AeroGen [tang2025aerogen], CC-Diff [zhang2024cc]), and natural images (LayoutDiffusion [zheng2023layoutdiffusion], GLIGEN [li2023gligen]). For a fair comparison, all models are re-trained using our dataset settings, following their official training details respectively.
  • Generation Fidelity. We use FID [heusel2017gans] and KID [binkowski2018demystifying] to assess perceptual quality, along with CMMD [jayasumana2024rethinking], which measures CLIP feature distances between generated and real images to evaluate layout alignment.
  • Layout Consistency. We report CAS [ravuri2019classification] using a pretrained classifier to assess object recognizability, and YOLOScore by applying a pretrained Oriented R-CNN [Xie_2021_ICCV] (w/. Swin Transformer backbone [liu2021swin], MMRotate) to generated images for instance-level consistency.
  • Shape Fidelity. To assess the geometric quality of generated instances, we perform pairwise comparisons with ground-truth shapes. Each instance pair is cropped, resized to $\times64$, and converted to edge maps. We compute five metrics: IoU, Dice, Chamfer Distance (CD), Hausdorff Distance (HD), and SSIM [wang2004image].
[…]
Figure 4: Qualitative results on DIOR, DOTA and HRSC2016. OF-Diff is more realistic and fidelity compared to other methods.
4.2 Qualitative Results
Diversity Results. The images generated by OF-Diff consistently present plausible textures and realistic object shapes, as shown in Figure 6 in Appendix. For instance, airplanes rendered at different orientations maintain coherent semantic relationships with their surrounding environments. Even in small-object scenes (some of which are grayscale remote-sensing images from the DOTA dataset), OF-Diff can still generate visually faithful and geometrically accurate results.
Table 1: Quantitative comparison with SOTA methods on DIOR and DOTA. We evaluate performance on generation fidelity (FID, KID, CMMD), layout consistency (CAS, YOLOScore) and trainability (mAP). OF-Diff demonstrates superior overall performance.